# Optimal Sparsity of Mixture-of-Experts Language Models for Reasoning Tasks

### v1

@@ 3.3 Optimal Sparsity for Iso-FLOP Budgets
In Figure 5, we plot the task-specific optimal sparsity (i.e. 1-TopK/Experts) against model performance under a fixed FLOPs budget. For QA benchmarks such as TriviaQA and HellaSwag, lower density (higher sparsity) consistently yields lower task loss and higher accuracy. This pattern aligns with prior studies showing that, when FLOPs are fixed to be constant, sparse models outperform denser models on QA tasks (Abnar et al., 2025). By contrast, on mathematical-reasoning benchmarks such as GSM8K and GSM-Plus, denser models outperform their sparser counterparts. At lower FLOPs, increasing sparsity still reduces loss and improves accuracy; however, once the FLOPs budget grows, denser models begin to perform better, achieving both lower loss and higher accuracy. This shift indicates that the optimal model density for reasoning tasks depends on compute budget: when a lot of FLOPs are available, denser models may be preferable.
@@ 3.4 Tokens per parameter
The Chinchilla scaling law (Hoffmann et al., 2022) establishes that, under a fixed compute budget, the optimal trade-off between model parameters and training tokens corresponds to approximately 20 tokens per parameter (TPP) for dense models. More recently, Roberts et al. (2025) refined this view by showing that the optimal TPP ratio is task-dependent: knowledge-based QA tasks, which are closer to memorization, benefit from lower TPP (i.e., more parameters), whereas reasoning-heavy tasks such as code generation benefit from higher TPP (i.e., more data). These findings highlight that TPP should be interpreted not as a universal constant, but as a task-sensitive scaling variable.
In our study, although we varied the number of experts while keeping the total FLOPs fixed, this implicitly altered the TPP measured with respect to total parameters. As shown in Figure 7, this variation reveals distinct behaviors across task categories. For memorization-oriented tasks such as TriviaQA and HellaSwag, performance improves monotonically as TPP decreases, consistent with the “parameter-hungry” characterization reported by Roberts et al. (2025). For reasoning tasks such as GSM8K and GSM-Plus, we observe a non-monotonic trend: accuracy peaks near TPP $\approx$ 20 and then declines, suggesting that excessively low TPP (i.e., too many parameters relative to tokens) can hurt reasoning performance.
Furthermore, our experiments reveal that active compute operationalized through the number of top-$k$ experts interacts strongly with TPP. Even at fixed TPP, models with larger top-$k$ values consistently outperform those with smaller top-$k$ on reasoning tasks. This indicates that, in MoE models, reasoning ability depends not only on the Total TPP but also on the balance between total and active parameters. In other words, the discussion of compute-optimal scaling in MoE architectures must explicitly consider both total parameter count and the number of activated parameters per token.
Figure 7: Effect of TPP on performance across different tasks. For TriviaQA and HellaSwag, performance improves as the number of parameters increases. In contrast, for reasoning-intensive tasks such as GSM8K and GSM-Plus, performance deteriorates when the number of parameters becomes too large, indicating that there exists an optimal data to total parameter ratio for these tasks. Even at fixed TPP, models with larger top-$k$ values consistently outperform those with smaller top-$k$ on reasoning tasks.
@@ 3.5 Impact of TTC and Post-Training on Downstream Performance
Test-Time Compute and RL post-training are standard for boosting reasoning on tasks such as mathematical problem solving. We therefore investigated whether performance declines reported above persist when applying (a) Test-Time Compute (TTC) and (b) RL post-training (GRPO). In Test-Time Compute, we evaluated GSM8K(Cobbe et al., 2021) in a purely zero-shot setting using Self-Consistency (SC) decoding(Wang et al., 2023), generating $2^{7}$ independent continuations per problem and selecting the most frequent answer. In Post-Training, we fine-tuned each model on the GSM8K training dataset using the GRPO algorithm (Shao et al., 2024). We followed the settings of Zhao et al. (2025) including reward function and fixed the learning rate constant across all model configurations.
@@
3.6 Influence of Optimization Hyperparameter
Figure 8: For reasoning tasks like GSM8K and GSM-Plus, the relationship between training loss and downstream performance is dependent on the choice of optimization hyperparameters. The learning rate also impacts downstream accuracy. For the maximum eigenvalue, we evaluated the maximum eigenvalue of fisher information matrix under a K-FAC approximation (Martens & Grosse, 2015; Eschenhagen et al., 2023). Following (Grosse et al., 2023), we calculate the maximum eigenvalues only for linear layers. We find that higher learning rates lead to a lower maximum eigenvalue, which is consistent with existing research indicating that convergence to flatter minima improves generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Jiang et al., 2020).
Figure 8 illustrates our empirical findings, obtained using a MoE architecture with 16 experts. By varying the learning rate, we evaluate performance on both QA benchmarks (TriviaQA, HellaSwag) and reasoning benchmarks (GSM8K, GSM-Plus). While QA benchmark performance like TriviaQA and HellaSwag remains largely invariant to these hyperparameters, reasoning benchmark performance like GSM8K and GSM-Plus are sensitive to the learning rates: when models converge to the same training loss, trainings with lower learning rates and smaller initialization scales yield superior downstream accuracy. These observations carry an important implication. Studies on generalization in large‐scale language models should incorporate rigorous reasoning benchmarks (such as GSM8K and GSM-Plus) rather than relying solely on validation loss curves or standard QA tasks to fully capture the impact of optimization‐induced implicit biases. This enables a more precise analysis on the generalization of LLMs.
3.7 Ablation on Depth
We conducted additional experiments using a 32 layer architecture. Motivated by prior reports suggesting that increased depth can improve performance (Liu et al., 2024b; Team, 2024; Ye et al., 2025), we evaluated whether deeper models exhibit similar trends in our setting. For the 32 layer configuration, we observed that the results align with the patterns discussed in the previous section, when analyzed through the lens of TPP, the behavior remains consistent with our earlier findings.
Figure 9: Effect of model depth on TPP-performance trade-offs.
@@ 3.8 Coding Task Ablations
We evaluate whether the sparsity performance trade offs observed for mathematical reasoning transfer to code generation. Unless otherwise noted, we reuse the same architecture and optimization hyperparameters as in the experimental setup. Models are trained on a 125B token corpus composed of 95B tokens from Stack-Edu Python (Allal et al., 2025) (high-quality educational Python code trained for four epochs following Muennighoff et al. (2023)) and 30B tokens from DCLM-dedup web text (Zyphra, 2024). We assess pass@1 accuracy on HumanEval (Chen et al., 2021) and MBPP  (Austin et al., 2021). See Appendix 3 for further details.
@@ 4 Discussion and Limitations
Dataset
All models are trained on a 125B-token corpus. This corpus is Chinchilla-optimal for dense models of comparable activated size (Hoffmann et al., 2022), yet two orders of magnitude smaller than the multi-trillion-token budgets now common for state-of-the-art MoE LLMs (DeepSeek-AI, 2025b; Qwen Team, 2025). Recent large models such as OLMo-2 and Qwen-3 adopt a multi-stage curriculum training, general web pre-training followed by mid-training on math and CoT data (OLMo, 2025; Qwen Team, 2025); we avoid this design to keep a fixed data distribution and a clean link between pre-train loss and downstream accuracy, but exploring staged curricula remains important future work. These caveats render our conclusions suggestive rather than prescriptive and motivate verification at trillion-token scale with richer reasoning corpora.
Model
We build on the Mixtral backbone and adopt the fine-grained expert segmentation of DeepSeek-MoE: each feed-forward block is split into $g=2$, so the effective expert count becomes $E\times g$ while the total parameter budget stays fixed. In conjunction with standard top-$k$ routing strategy ($k\!\in\!\{2,4,8,16\}$) and the auxiliary importance / load-balance loss of Shazeer et al. (2017), our hyperparameter sweep evaluates configurations with up to $256$ active experts. This is contrast to contemporary MoE variants such as Qwen‑3, which primarily differ from Mixtral by the integration of only QK‑Norm and a global load‑balancing regularizer. These modifications are negligible in comparison to the scale of changes evaluated in our experiments. For scaling, the number of experts is more influential than minor structural details. We explore configurations with up to 256 experts and a top-16 routing strategy, which offers a sufficiently broad range for our purposes. We acknowledge that gating design choices, such as the formulation of the load-balance loss, might affect how expert scaling influences performance; we leave this for future work. The patterns we report are intended as provisional observations rather than definitive rules. We encourage further studies to examine these effects at larger model scales.
@@ 5 Conclusion
In this paper, we investigated the optimal sparsity of MoE language models through the lens of downstream task performance. By training families of Mixtral-style MoEs with various number of experts, top-k routing, and model width, and by evaluating them across pre-training, GRPO post-training, and test-time compute, we show that the classical “more experts is better” rule holds for knowledge-oriented benchmarks such as TriviaQA and HellaSwag, but not for mathematical reasoning benchmarks. On reasoning tasks, downstream task loss starts to rise, and accuracy to fall, once total parameters grow at a certain point; in this regime, models with more active parameters may achieve lower optimal task loss, whereas those with extreme sparsity over-fit despite lower pre-training loss. Neither reinforcement-learning post-training nor additional test-time compute removes this trade-off. These findings update current scaling practice. When computational budget is fixed, allocating FLOPs to extra experts improves memorization, but improving reasoning ability requires matching growth in active parameters or even shifting toward denser MoE layers once enough compute is available.
@@ Acknowledgements
This work was supported by the “R&D Hub Aimed at Ensuring Transparency and Reliability of Generative AI Models” project of the Ministry of Education, Culture, Sports, Science and Technology. We used ABCI 3.0 provided by AIST and AIST Solutions with support from “ABCI 3.0 Development Acceleration Use. This work used computational resources TSUBAME4.0 supercomputer provided by Institute of Science Tokyo through the HPCI System Research Project (Project ID: hp240170).
@@ C.2 Test-Time Compute
To set up Test Time Compute appropriately, we investigate how varying the number of prompt shots affected each expert’s behavior (Figure 13). Few shot performance is unstable and dropped significantly for models with a small number of experts, so we use zero shot inference for Test Time Compute. When few shot chain of thought is used to standardize answer formats, the provided demonstration steps can be internalized as a fixed reasoning pattern by the model. As a result, the model’s inherent inference capabilities may not be fully expressed, and its ability to generalize to novel problems could be hindered (Kojima et al., 2022).
@@ C.3 GSM8K Overfitting Analysis
The results are shown in Figure 17 and 18. We observe that the trends with respect to sparsity on GSM8K remain unchanged, both for Pass@1 and TTC metrics. This indicates that while GSM8K training data and its synthetic derivatives do improve GSM8K scores, they do not alter the underlying performance trends. However, after post-training, we observe some changes in these trends, which we leave as future work to investigate further.
@@ C.5 Detailed Results for Coding Tasks
This appendix provides detailed results for the coding task ablations, mirroring the analyses for mathematical reasoning presented in the main text. Specifically, we detail the non-monotonic relationship between scaling and downstream performance, showing how both task loss (Figure 20, 21) and accuracy (Figure 22) can degrade as pre-training loss improves. We then analyze key architectural factors, including the impact of MoE sparsity (Figure 23) and the optimal Tokens-per-Parameter (TPP) ratio (Figure 24).

### v2

@@ 3.3 Optimal Sparsity for Iso-FLOP Budgets
In Figure 5, we show the task-specific optimal sparsity (i.e. 1-TopK/Experts) against model performance under a fixed FLOPs budget. For memorization benchmarks, lower density (higher sparsity) consistently yields lower task loss and higher accuracy. This pattern aligns with prior studies showing that, when FLOPs are fixed to be constant, sparse models outperform denser models on QA tasks (Abnar et al., 2025). By contrast, on mathematical-reasoning benchmarks, denser models outperform their sparser counterparts. At lower FLOPs, increasing sparsity still reduces loss and improves accuracy; however, once the FLOPs budget grows, denser models begin to perform better, achieving both lower loss and higher accuracy. This shift indicates that reasoning skills admit a compute-dependent optimum, rather than monotonically favoring either sparsity or density.
@@ 3.4 Tokens per parameter
The Chinchilla scaling law (Hoffmann et al., 2022) establishes that, under a fixed compute budget, the optimal trade-off between model parameters and training tokens corresponds to approximately 20 tokens per parameter (TPP) for dense models. More recently, Roberts et al. (2025) refined this view by showing that the optimal TPP ratio is task-dependent: memorization skills benefit from lower TPP (i.e., more parameters), whereas reasoning skills benefit from higher TPP (i.e., more data). These findings highlight that TPP should be interpreted not as a universal constant, but as a task-sensitive scaling variable.
[…]
Furthermore, our experiments reveal that active compute operationalized through the number of top-$k$ experts interacts strongly with TPP. Even at fixed TPP, models with larger top-$k$ values consistently outperform those with smaller top-$k$ on reasoning tasks. This indicates that, in MoE models, reasoning ability depends not only on the Total TPP but also on the balance between total and active parameters. In other words, the discussion of compute-optimal scaling in MoE architectures must explicitly consider both total parameter count and the number of activated parameters per token. We further note that depth ablations (Appendix C.6) exhibit the same non-monotonic dependence on TPP.
Figure 7: Effect of TPP on performance across different tasks. For TriviaQA and HellaSwag, performance improves as the number of parameters increases. In contrast, for reasoning skills, performance deteriorates when the number of parameters becomes too large, indicating that there exists an optimal total tokens per paramete ratio for these tasks. Even at fixed TPP, models with larger top-$k$ values consistently outperform those with smaller top-$k$ on reasoning tasks.
@@ 3.5 Impact of TTC and Post-Training on Downstream Performance
Test-Time Compute and RL post-training are standard for boosting reasoning on tasks such as mathematical problem solving. We therefore investigated whether the performance trade-offs observed above persist or shift when applying (a) Test-Time Compute (TTC) and (b) RL post-training (GRPO). In Test-Time Compute, we evaluated GSM8K(Cobbe et al., 2021) in a zero-shot setting using Self-Consistency (SC) decoding(Wang et al., 2023), generating $2^{7}$ independent continuations per problem and selecting the most frequent answer. In Post-Training, we fine-tuned each model on the GSM8K training dataset using the GRPO algorithm (Shao et al., 2024). We followed the settings of Zhao et al. (2025) including reward function and fixed the learning rate constant across all model configurations.
@@ 3.6 Coding Task Ablations
We evaluate whether the sparsity performance trade offs observed for mathematical reasoning transfer to code generation. Unless otherwise noted, we reuse the same architecture and optimization hyperparameters as in the experimental setup. Models are trained on a 125B token corpus composed of 95B tokens from Stack-Edu Python (Allal et al., 2025) (high-quality educational Python code trained for four epochs following Muennighoff et al. (2023)) and 30B tokens from DCLM-dedup web text (Zyphra, 2024). We assess pass@1 accuracy on HumanEval (Chen et al., 2021) and MBPP  (Austin et al., 2021). See Table 3 in Appendix for further details.
@@ 4 Discussion and Limitations
All models are trained on a 125B token corpus, which is Chinchilla-optimal for dense models of comparable activated size (Hoffmann et al., 2022). Except for the widest $d=2048$, top-16 setting, our runs are in or near the Chinchilla-optimal regime, or fall into the overtraining regime. Recent frontier models typically train with a fixed token budget (Grattafiori et al., 2024; Qwen Team, 2025), so holding tokens fixed in our experiments is appropriate. Nevertheless, with larger training corpora, one could train larger models at higher tokens-per-parameter, potentially shifting the optimal sparsity toward sparser configurations, even for reasoning skills. We leave this as a direction for future work.
Our study does not attempt to exhaustively explore all architectural and training settings. The design space of MoE models is vast and it is infeasible to cover every possible combination. Instead, we focused on a representative but systematic sweep over width, expert count, and top-$k$ routing, which already reveals new regularities that prior work did not capture. In particular, we identify an inverted-U relationship between sparsity and reasoning performance that contradicts the monotonic trends often assumed in scaling analyses. Our analysis further shows that this effect can be parsimoniously explained in terms of TPP and active FLOPs: reasoning performance peaks near a task-specific TPP optimum and degrades when models are either too parameter-heavy or too data-starved, even as training loss continues to fall. Existing scaling laws are incomplete and require refinement.
@@ 5 Conclusion
We investigated the optimal sparsity of MoE language models through a large-scale exploration of Mixtral-style architectures, varying expert count, top-$k$ routing, and width, and evaluating across pre-training, RL post-training, and test-time compute. Our results reveal two central insights. First, FLOPs matter: downstream reasoning quality is determined not by pre-training loss alone, but by the number of active FLOPs at both train and test time. Larger top-$k$ consistently outperforms smaller ones even when pre-training loss is matched. Second, TPP matters: the tokens-per-parameter ratio governs task-specific scaling. Memorization tasks are parameter-hungry and benefit from sparsity, while reasoning tasks are data-hungry and peak near $20$ tokens per parameter, degrading when TPP becomes too low. Together, these findings refine current scaling practice. Sparsity and more experts improve memorization under fixed budgets, but reasoning requires balancing active FLOPs with TPP. In high-compute regimes, the optimal density depends jointly on compute and data: with limited data, denser MoE layers preserve reasoning, while with abundant data, greater sparsity remains advantageous.
@@ Acknowledgements
[…]
@@ C.2 Test-Time Compute
[…]
@@ C.3 GSM8K Overfitting Analysis
The results are shown in Figure 15 and 16. We observe that the trends with respect to sparsity on GSM8K remain unchanged, both for Pass@1 and TTC metrics. This indicates that while GSM8K training data and its synthetic derivatives do improve GSM8K scores, they do not alter the underlying performance trends. However, after post-training, we observe some changes in these trends, which we leave as future work to investigate further.
@@ C.5 Detailed Results for Coding Tasks
This appendix provides detailed results for the coding task ablations, mirroring the analyses for mathematical reasoning presented in the main text. Specifically, we detail the non-monotonic relationship between scaling and downstream performance, showing how both task loss (Figure 18, 19) and accuracy (Figure 20) can degrade as pre-training loss improves. We then analyze key architectural factors, including the impact of MoE sparsity (Figure 21) and the optimal Tokens-per-Parameter (TPP) ratio (Figure 22).
@@
C.6 Ablation on Depth
We conducted additional experiments using a 32 layer architecture. Motivated by prior reports suggesting that increased depth can improve performance (Liu et al., 2024; Team, 2024; Ye et al., 2025), we evaluated whether deeper models exhibit similar trends in our setting. For the 32 layer configuration, we observed that the results align with the patterns discussed in the previous section, when analyzed through the lens of TPP, the behavior remains consistent with our earlier findings.
Figure 23: Effect of model depth on TPP-performance trade-offs.
C.7 Influence of Optimization Hyperparameter
Figure 24: For reasoning skills, the relationship between training loss and downstream performance is dependent on the choice of optimization hyperparameters. The learning rate also impacts downstream accuracy. For the maximum eigenvalue, we evaluated the maximum eigenvalue of fisher information matrix under a K-FAC approximation (Martens & Grosse, 2015; Eschenhagen et al., 2023). Following (Grosse et al., 2023), we calculate the maximum eigenvalues only for linear layers. We find that higher learning rates lead to a lower maximum eigenvalue, which is consistent with existing research indicating that convergence to flatter minima improves generalization (Hochreiter & Schmidhuber, 1997; Keskar et al., 2017; Jiang et al., 2020).
Figure 24 illustrates our empirical findings, obtained using a MoE architecture with 16 experts. While memorization skills remains largely invariant to these hyperparameters, reasoning skills are sensitive to the learning rates: when models converge to the same training loss, trainings with lower learning rates and smaller initialization scales yield superior downstream accuracy. These observations carry an important implication. Studies on generalization in large‐scale language models should incorporate rigorous reasoning benchmarks rather than relying solely on validation loss curves or standard QA tasks to fully capture the impact of optimization‐induced implicit biases. This enables a more precise analysis on the generalization of LLMs.
[…]